OpenStudy (anonymous):

what is the standard deviation of 623658?

OpenStudy (amistre64):

you might want to provide more context, or a better description of what it is you are talking about

OpenStudy (amistre64):

is that a number? or a set of numbers that have no commas? or what?

OpenStudy (anonymous):

its a set of numbers with no commas

OpenStudy (amistre64):

lol, commas would be nice

OpenStudy (amistre64):

find the mean subtract the mean from each datapoint square the results divide by how many elements there are then take the sqrt

OpenStudy (amistre64):

\[V(x)=E(x-\mu)^2\] \[V(x)=E(x^2-2\mu x+\mu^2)\] \[V(x)=E(x^2)-2\mu ~E(x)+\mu^2,~but~ E(x)=\mu\] \[V(x)=E(x^2)-2\mu^2+\mu^2\] \[V(x)=E(x^2)-\mu^2\]